OpenStudy (anonymous):

Mariana had a $20 gift card to an online music store. She spent the entire amount on songs, which cost $1 each,and music videos, which cost $2 each. Mariana bought 5 more songs than music videos. write and solve a system of equations to find the number of songs and the number of videos Mariana bought.

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (triciaal):

20 = S + 2V S = V + 5 s = song and v = video

OpenStudy (anonymous):